OpenStudy (anonymous):

Find the coordinates of the stationary point of the curve y=ln(x²-6x+10).

OpenStudy (anonymous):

all stat pts-> dy/dx = 0

OpenStudy (anonymous):

i.e. differentiate your function and solve

OpenStudy (anonymous):

To differentiate log e, Take whatever is in the bracket, put it on the bottom of a fraction (denominator), differentiate the bottom of the fraction then place the differentiated expression on the top of the fraction (numerator)

OpenStudy (anonymous):

\[\frac{ dy }{ dx } \ln(x ^{2} - 6x +10) = \frac{ 1 }{ x ^{2} - 6x + 10 }\]

OpenStudy (anonymous):

@Skaematik ?

OpenStudy (phoenixfire):

@curiousshubham Almost. You need to use the chain rule to differentiate this.\[{d \over dx}(f(g(x)))=f^{\prime}(g(x))g^{\prime}\]\[{d \over dx}\ln(x^2-6x+10)\] Let \[f(x)=ln(x)\] and \[g(x)=x^2-6x+10\] Find \[f^{\prime}={d \over dx}ln(x)\]\[g^{\prime}={d \over dx}(x^2-6x+10)\] And substitute into the equation up the top.

OpenStudy (anonymous):

@PhoenixFire I got \[\frac{ 1 }{ x ^{2} - 6x +10 }\]

OpenStudy (phoenixfire):

that's df/dx. you need to multiply by the derivative of the inside function (dg/dx)

OpenStudy (anonymous):

y = ln(f(x)) y' = f'(x)/f(x) , right?

OpenStudy (phoenixfire):

yup.

OpenStudy (phoenixfire):

So you get \[{d \over dx}\ln(x^2-6x+10)={{2x-6} \over {x^2-6x+10}}\]

OpenStudy (anonymous):

yup!

OpenStudy (anonymous):

@PhoenixFire Equatting it with 0?

OpenStudy (phoenixfire):

Yeah, equate to 0 and solve for x. for each value of x plug it into the original equation to get your y. then you will have your coordinates.